OpenStudy (anonymous):

I WILL FAN AND MEDAL IF YOU CAN HELP ME WITH SOME SIMLE ALGEBRA!!!!!! :) Simplify the rational expression by rationalizing the denominator. equation below

OpenStudy (anonymous):

\[\frac{ 4 \sqrt{150} }{ \sqrt{189x} }\]

OpenStudy (cggurumanjunath):

150=2*75=2*3*25=2*3*5*5

OpenStudy (cggurumanjunath):

4*sqrt (150) =4*5*sqrt (2*3)=20*sqrt 6

OpenStudy (anonymous):

Wait which part do you want me to solve?

OpenStudy (cggurumanjunath):

189=3*63=3*21*3=3*3*3*7

OpenStudy (anonymous):

20*sqrt6 =?

OpenStudy (cggurumanjunath):

sqrt (189)=sqrt(3*3*3*7)=3*sqrt (21)

OpenStudy (anonymous):

@CGGURUMANJUNATH no offense but I think your confusing @Darealest123 lol

OpenStudy (cggurumanjunath):

numerator = 20*sqrt (6)

OpenStudy (anonymous):

Also answer choices are: a 20*sqrt150/sqrt189x b 4/sqrt189x c 4*sqrt39x/189x d 16/189x

OpenStudy (cggurumanjunath):

denominator = sqrt (189*x) = 3*sqrt (21*x)

OpenStudy (cggurumanjunath):

n/d

OpenStudy (anonymous):

Darealest what grade is this?

OpenStudy (anonymous):

grade? you mean course?

OpenStudy (cggurumanjunath):

numerator =n denominator =d

OpenStudy (anonymous):

no like what grade math is this lol I know its algebra but like what level? 10th, 9th, 11th

OpenStudy (cggurumanjunath):

20*sqrt (6) / 3 *sqrt (21x)

OpenStudy (anonymous):

so it would be 20*sqrt6/3*sqrt21*x

OpenStudy (anonymous):

but thats not an answer choice so do i simplify the numbers?

OpenStudy (cggurumanjunath):

yes !

OpenStudy (anonymous):

48.9/?

OpenStudy (anonymous):

right so far or wrong